OpenStudy (anonymous):

factor completely. cd+2c+8d+16

OpenStudy (swaugh):

answer choices

OpenStudy (swaugh):

if they are no answer choices try (c+8)(d+2)

OpenStudy (anonymous):

\[cd+2c(d+2)\\ 8d+16=8(d+2)\]

OpenStudy (anonymous):

so you have \[2c(d+2)+8(d+2)=(2c+8)(d+2)\]